Financial data for Joel de Paris, Inc., for last year follow: Joel de Paris, Inc. Balance Sheet Beginning Balance Ending Balance Assets Cash $ 131,000 $ 127,000 Accounts receivable 340,000 486,000 Inventory 565,000 482,000 Plant and equipment, net 844,000 825,000 Investment in Buisson, S.A. 396,000 434,000 Land (undeveloped) 248,000 254,000 Total assets $ 2,524,000 $ 2,608,000 Liabilities and Stockholders' Equity Accounts payable $ 389,000 $ 338,000 Long-term debt 981,000 981,000 Stockholders' equity 1,154,000 1,289,000 Total liabilities and stockholders' equity $ 2,524,000 $ 2,608,000 Joel de Paris, Inc. Income Statement Sales $ 3,990,000 Operating expenses 3,351,600 Net operating income 638,400 Interest and taxes: Interest expense $ 122,000 Tax expense 203,000 325,000 Net income $ 313,400 The company paid dividends of $178,400 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%. Required: 1. Compute the company's average operating assets for last year. 2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.) 3. What was the company’s residual income last year?
Answer #1
1
Beginning Operating assets 1880000 =2524000-396000-248000
Ending Operating assets 1920000 =2608000-434000-254000
Average Operating assets 1900000 =(1880000+1920000)/2
2
Margin = Net operating income/Sales
Turnover = Sales/Average Operating assets
ROI = Margin*Turnover
Margin 16.00% =638400/3990000
Turnover 2.10 =3990000/1900000
ROI 33.60% =16%*2.1
3
Net operating income 638400
Less: Minimum required return 285000 =1900000*15%
Residual Income 353400